# Quantum Enhanced Anomaly Detection for ADS-B Data using Hybrid Deep Learning

### v1

@@ VI Discussion
We emphasize that the models in this study were developed with a focus on exploring the impact of quantum properties on anomaly detection on ADS-B data and not with the intention of creating the most effective models. The results from this study indicate that QML can effectively process ADS-B data for anomaly detection. The performance of the quantum models sometimes outperformed classical models, but it also performed slightly worse in other cases, indicating that hybrid models still require exploration and to establish an advantage over classical models. However, the models showed similar tendency and bias regarding the different techniques used. These important findings suggest that, when quantum devices will reach the Fault-Tolerant Quantum Computer (FTQC) era, the aeronautic domain could benefit from quantum properties to enhance its cybersecurity systems. Machine learning represents a prime candidate for quantum speedups as even polynomial improvements can yield substantial benefits due to the high computational demands of data processing [16]. As every second counts in aviation cybersecurity, the application of QML on the ADS-B is an excellent candidate.
It is also important to consider that all experiments were conducted on noise-free quantum simulators. The noise-free simulator from Pennylane does not account for real-world noise, decoherence, and gate errors inherent in quantum hardware. Moreover, running experiments on real quantum computers remains costly, which limits the feasibility of large-scale testing. While the impact of noise on current quantum computers is significant, the technology is continually advancing and new developments are improving hardware reliability and scalability. Relying on the IBM roadmap for quantum technologies, it is predicted that in the next 5 years, quantum computers will be more reliable in that sense [31]. Accordingly, this work focused on the employment of QML, rather than on the impacts of imperfections on the quality of results. Beyond these general limitations, specific challenges arise from the nature of the project itself. The aerospace industry demands highly deterministic and reliable systems. Further exploration is needed to achieve reliable performance to align well with the strict safety requirements of aviation.

### v2

@@ VI Discussion
We emphasize that the models in this study were developed with a focus on exploring the impact of quantum properties on anomaly detection on ADS-B data and not with the intention of creating the most effective models. The results from this study indicate that QML can effectively process ADS-B data for anomaly detection. The performance of the quantum models sometimes outperformed classical models, but it also performed slightly worse in other cases, indicating that hybrid models still require exploration and to establish an advantage over classical models. However, the models showed similar tendency and bias regarding the different techniques used. These important findings suggest that, when quantum devices will reach the Fault-Tolerant Quantum Computer (FTQC) era, the aeronautic domain could benefit from quantum properties to enhance its cybersecurity systems. Machine learning represents a prime candidate for quantum speedups as even polynomial improvements can yield substantial benefits due to the high computational demands of data processing [Schuld2021]. As every second counts in aviation cybersecurity, the application of QML on the ADS-B is an excellent candidate.
It is also important to consider that all experiments were conducted on noise-free quantum simulators. The noise-free simulator from Pennylane does not account for real-world noise, decoherence, and gate errors inherent in quantum hardware. Moreover, running experiments on real quantum computers remains costly, which limits the feasibility of large-scale testing. While the impact of noise on current quantum computers is significant, the technology is continually advancing and new developments are improving hardware reliability and scalability. Relying on the IBM roadmap for quantum technologies, it is predicted that in the next 5 years, quantum computers will be more reliable in that sense [ibm2024roadmap]. Accordingly, this work focused on the employment of QML, rather than on the impacts of imperfections on the quality of results. Beyond these general limitations, specific challenges arise from the nature of the project itself. The aerospace industry demands highly deterministic and reliable systems. Further exploration is needed to achieve reliable performance to align well with the strict safety requirements of aviation.